OpenStudy (anonymous):

what is the equation in point-slope form of the line passing through (1,9) and (-1,11)

OpenStudy (anonymous):

y = -x +10

OpenStudy (anonymous):

\[M = (y_2 - y_1)/(x_2 - x_1)\]

OpenStudy (anonymous):

= 11 - 9 / -1 - 1 = 2/ -2 m = -1

OpenStudy (anonymous):

\[y - y_1 = m(x - x_1)\]

OpenStudy (anonymous):

y - 9 = -1(x - 1) y - 9 = -x +1 y = -x +1 + 9 y = -x + 10